In 28 years, a woman will be twice as old as she will be in two years. How old is she now?
Answer by Cromlix(4381)

Woman's age = x
(x + 28) = 2(x + 2)
x + 28 = 2x + 4
Collext like terms;
x - 2x = 4 - 28
-x = -24
x = 24.
The woman is 24 years old at present.
Hope this helps.
;-)